# Nuclear matter properties from chiral-scale effective theory including a dilatonic scalar meson

### v1

@@ I Introduction
Study on nuclear matter (NM) has long been crucial for understanding both nuclear force and neutron star (NS) structures (see, e.g., Refs. Serot and Walecka (1986); Brown and Rho (2002); Fukushima and Hatsuda (2011); Lattimer and Prakash (2016); Baym et al. (2018); Ma and Rho (2020a); Brandes and Weise (2024) and references therein). The properties of NM are highly sensitive to details of nucleon interactions. A popular and widely used approach to describe these interactions is the one-boson-exchange (OBE) model and its variants, such as the Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989). These models typically involve $\pi$, $\sigma$, $\omega$ and $\rho$ meson exchanges, covering the effective range of nucleon forces from $0.5\ \mathrm{fm}$ to $2\ \mathrm{fm}$. Utilizing these models, NM properties can be calculated using the relativistic mean field (RMF) approach Walecka (1974); Wiringa et al. (1988), which is the most practical and economical framework to introduce density effects.
@@
It is recognized that Walecka-type models lack considerations of QCD symmetry patterns, the valid region of effective operators, and theoretical errors. Chiral effective field theory ($\rm{\chi EFT}$), thanks to the pioneering works by Weinberg Weinberg (1979, 1991), offers a powerful framework for studying nuclear forces at long ranges, anchored on QCD symmetry. To develop a realistic model of nuclear forces, it is accepted that vector mesons—$\rho$ and $\omega$— and isoscalar-scalar meson $\sigma$ are indispensable. The former can be regarded as gauge fields of hidden local symmetry (HLS) in $\rm{\chi EFT}$  Bando et al. (1985, 1988); Harada and Yamawaki (2003). However, the inclusion of the $\sigma$ meson as an independent degree of freedom in $\rm{\chi EFT}$ is not straightforward, since the fourth scalar component of the chiral four-vector is integrated out when transitioning from the linear sigma model to the nonlinear sigma model—the leading term of $\rm{\chi EFT}$. Considering that the lowest-lying scalar meson $\sigma$ has a mass roughly equivalent to that of the kaon, which is a Nambu-Goldstone (NG) boson of chiral symmetry breaking in three-flavor $\rm{\chi EFT}$ ($\chi\rm EFT_{3}$), by supposing that QCD has a nonperturbative infrared fixed point, Crewther and Tunstall Crewther and Tunstall (2015); Catà et al. (2019) proposed that the lowest-lying scalar meson can be considered as the NG boson (dilaton) of the spontaneous breaking of scale symmetry. Consequently, $\rm{\chi EFT}$ is extended to include the scalar meson, resulting in the scale-chiral effective theory ($\chi\rm EFT_{\sigma}$).
@@
Based on the terminologies of $\chi\rm EFT_{\sigma}$ and HLS, the $\chi\rm EFT$ with baryons is constructed in Refs. Lee et al. (2015); Li et al. (2017) at the leading chiral order to discuss nucleon interactions. It is denoted as bsHLS, with ‘b’ for baryon and ‘s’ for scale. In bsHLS, the potential from meson exchanges in OBE models can be reproduced by expanding bsHLS to the first order of linear field couplings but with additional symmetry considerations. By using the low-momentum potential $V_{\rm lowk}$ renormalization group approach Bogner et al. (2003); Holt et al. (2007), taking the “leading order scale symmetry (LOSS)” approximation where the trace anomaly effect enters only through the dilaton potential, which breaks scale symmetry explicitly and spontaneously, we found that the chiral-scale EFT with few parameters can successfully describe not only NM at the saturation density but also the compact-star matter at $n=(5-7)n_{0}$ Ma and Rho (2020a). A novel phenomena, which was not realized, is that in nuclear matter at densities $n=(2-4)n_{0}$ the sound velocity of NM saturates the conformal limit $v_{s}^{2}=1/3$, but the trace of the energy-momentum tensor does not vanish, that is, the NM exhibits a pseudo-conformal structure Ma and Rho (2019a, b); Ma et al. (2019).
@@
In this work, we study NM properties using bsHLS with the RMF method. The corrections to LOSS in the baryonic part, which have been found crucial for understanding the quenched $g_{A}$ value in the super-allowed Gamow-Teller transitions of heavy nuclei Lubos et al. (2019); Ma and Rho (2020b). After identifying the NM properties around $n_{0}$, it was found that bsHLS using the RMF method can yield reasonable results. The data can be reproduced with the choice of $\beta^{\prime}=\frac{\partial\beta}{\partial\alpha}\sim 1$, consistent with the results of Refs. Ma and Rho (2018); Shao and Ma (2022), where a dilaton limit fixed point is assumed in the medium. The combination of $f_{\chi}m_{\sigma}$ is constrained to be $\simeq 2.3\times 10^{5}$ MeV2 which is consistent with what estimated before in the skyrmion crystal approach such that physically interesting results can be obtained Lee et al. (2003); Park et al. (2008). By comparing the symmetry energy $E_{\rm sym}$ and incompressibility coefficient $K(n)$ obtained from our bsHLS and Walecka-type models, we find that the bsHLS can make the symmetry energy stiff at subsaturation density but soft at intermediate densities to meet the constraints of GW170817 Abbott et al. (2018) and the neutron skin thickness of ${}^{208}\rm Pb$ Reed et al. (2021) simultaneously, without introducing additional freedoms such as the $\delta$ meson in Refs. Li et al. (2022); Ma and Ma (2024). In addition, we find that the incompressibility coefficient of bsHLS surges at intermediate densities, while the Walecka-type models exhibit gentler behaviors, resulting in a better description of NS structures using bsHLS. The maximum mass of NS can nearly reach $3M_{\odot}$ with a pure hadron phase, meeting the constraints of GW170817 Abbott et al. (2018, 2019) and PSR J0740+6620 Miller et al. (2021), whereas Walecka-type models can only reach $2M_{\odot}$ within these constraints, as analyzed in Ref. Guo et al. (2024). We find that this is due to the kink behavior of $\sigma$ expectations at intermediate densities, induced by the nonlinear realization of scale symmetry. Moreover, the value of $\beta^{\prime}$, behavior of $\langle\chi\rangle^{*}$ and Brown-Rho scaling (B-R scaling) at different densities can significantly affect NS structures, indicating a relationship between QCD symmetry patterns and macroscopic phenomena.
@@
In this section, we establish the theoretical framework of bsHLS and the EoS of NM from the RMF approach. Here, we only interest in NM made of nucleons, we reduce the bsHLS Lagrangian to the two-flavor case, $u$ and $d$. The leading order bsHLS Lagrangian $\mathcal{L}$ can be decomposed into the baryonic part $\mathcal{L}_{B}$ and the mesonic part $\mathcal{L}_{M}$:
@@
$g_{\omega NN}=\frac{1}{2}\left(g_{V_{\rho}}+g_{V_{0}}-1\right)g_{\omega}\ ,$
$g_{\rho NN}=\frac{1}{2}\left(g_{V_{\rho}}-1\right)g_{\rho}\ ,$
@@ III Phenomenological analysis
In the phenomenological analysis, we choose vacuum values $f_{\pi}=92.4~{}\rm MeV$, $m_{N}=939~{}\rm MeV$, $m_{\omega}=783~{}\rm MeV$ and $m_{\rho}=765~{}\rm MeV$ Workman et al. (2022). The free parameters are $M_{\sigma}=m_{\sigma}f_{\chi}$, $\beta^{\prime}$, $r$, $g_{\omega NN}$, $g_{\rho NN}$, $g^{SSB}_{\omega NN}$ and $g^{SSB}_{\rho NN}$, which can be estimated by the properties of NM around saturation density $n_{0}$. The calculated results of NM properties are listed in Table 1, and the corresponding low energy constants (LECs) are given in Table 2. It can be seen that NM properties obtained from both bsHLS-L and bsHLS-H are consistent with empirical values.
@@ III.1 Nuclear matter properties
We first study the NM properties. From Table 1, one can see that, with appropriate choices of parameters in Table 2, our bsHLS can yield NM properties around (sub)saturation density, satisfying the constraints from empirical data.
@@ III.2 Neutron star structures
Next, we study the NS structures using the EOSs discussed above. The results of NS structures are shown in Fig. 2 and Table 4.
Figure 2: The M-R relations from bsHLS and Walecka-type models. The constraints are estimated from Refs. Abbott et al. (2018, 2019); Riley et al. (2019, 2021)
Table 4: Tidal deformations from bsHLS and Walecka-type models defined in Eq. (14) and Table 3 for NS with mass of $1.4\ M_{\odot}$.
@@
From Fig. 2, one can see that the mass-radius (M-R) relation obtained from bsHLS-H falls within the constraints of astro-observations. Although the result from FSU-$\delta 6.7$ satisfies these constraints, it requires introducing a new degree of freedom $\delta$. The Walecka-type models without $\delta$ considered in this work cannot provide results within the constraints of NSs, though they reproduce the NM properties around $n_{0}$ Horowitz and Serot (1981); Reinhard et al. (1986); Sugahara and Toki (1994). This highlights the advantage of bsHLS in interpreting the structures of nuclei and NSs within a unified framework. It should be emphasized that the present analysis indicates that, in order to have realistic parameter space of nuclear model, including Walecka-type models, the data of both NM properties and NS structures should be considered in the statistical analysis simultaneously, as discussed in Ref. Guo et al. (2024). When meeting the NS constraints, the $M_{\rm max}$ of NSs predicted by these Walecka-type models is around $2M_{\odot}$, as shown in Ref. Özel and Freire (2016); Guo et al. (2024). This makes bsHLS more appealing, as the predicted $M_{\rm max}$ of bsHLS-H is $\sim 2.8M_{\odot}$. This predicted $M_{\rm max}$ may have profound implications for addressing the gap problem in the continuous mass distribution of supernova remnants Bailyn et al. (1998); Farr et al. (2011) and for understanding the gravitational wave event GW190814 Abbott et al. (2020).
@@
In summary, the above discussion indicates that bsHLS-H is a reasonable model for NM and NS. This implies that the scaling parameter $r\approx 0.19$, $\beta^{\prime}\approx 1.15$, $M_{\sigma}=m_{\sigma}f_{\chi}\approx 2.3\times 10^{5}~{}\rm MeV^{2}$ and the couplings between vector mesons and nucleons $g_{\rho NN}\approx 4.17$, $g_{\omega NN}\approx 11.0$. The scaling parameter $r\approx 0.19$ is consistent with pion-nuclei bound state data Kienle and Yamazaki (2004). The $\beta^{\prime}\approx 1.15$ agrees with estimates from skyrmion crystal approach Ma and Rho (2018); Shao and Ma (2022). If $f_{\chi}$ is taken to be $3f_{\pi}\approx 270~{}\rm MeV$, the power-counting mechanism of Crewther and Tuntall remains valid Crewther and Tunstall (2015), while $m_{\sigma}\approx 850~{}\rm MeV$. The coupling constant $g_{\rho NN}\approx 4.17$ aligns with the results from the OBE potential analysis of nucleon interactions Erkelenz (1974), and $g_{\omega NN}\approx 11.0$ is in agreement with the analyses of nucleon-nucleon scatterings Backman et al. (1985).
@@ III.3 The patterns of scale symmetry and phenomenologies
The difference between bsHLS-L and bsHLS-H becomes significant in the M-R relation, as shown in Fig. 2, though it is not distinguishable from the NM properties around saturation density. The key difference lies in $\beta^{\prime}$, which connects to the existence of the pseudo-conformal structure at high densities Shao and Ma (2022). The $\beta^{\prime}$ of bsHLS-H aligns with the constraints of the pseudo-conformal limit, whereas the $\beta^{\prime}$ of bsHLS-L does not. This indicates the influence of scale symmetry on NS structures.
@@
And, as shown in Fig. 3, the M-R relations are roughly consistent with observational constraints, but there is a decrease of $M_{\rm max}$ compared to bsHLS-H.
@@
Additionally, the NM properties were investigated with B-R scaling turned off. We find that with the parameter set bsHLS-N:$M_{\sigma}=1.11\times 10^{5}\ \rm MeV^{2}$, $\beta^{\prime}=1.10$, $g_{\omega NN}=-11.0$, $g_{\rho NN}=4.17$, $g^{SSB}_{\omega NN}=-192$ and $g^{SSB}_{\rho NN}=4.85$, the NM properties can still be reproduced (e.g., $n_{0}=0.16\ \rm{fm}^{-3}$, $e_{0}=-16.0\ \rm{MeV}$, $K_{0}=241\ \rm{MeV}$, $E_{{\rm sym}}(n_{0})=29.3\ \rm{MeV}$). However, this approach results in a much softer scale symmetry parameter $\langle\chi\rangle^{*}$ flow compared to the cases with B-R scaling. As a result, the M-R relation becomes unnatural and deviates from the constraints, as shown in Fig. 5.

### v3

@@ I Introduction
Study on nuclear matter (NM) has long been crucial for understanding both nuclear force and neutron star (NS) structures (see, e.g., Refs. Serot and Walecka (1986); Brown and Rho (2002); Fukushima and Hatsuda (2011); Lattimer and Prakash (2016); Baym et al. (2018); Ma and Rho (2020a); Brandes and Weise (2024) and references therein). The properties of NM are highly sensitive to details of nucleon interactions. A popular and widely used approach to describe these interactions is the one-boson-exchange (OBE) model and its variants, such as the Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989). These models typically involve $\pi$, $\sigma$, $\omega$ and $\rho$ meson exchanges, covering the effective range of nucleon forces from $0.5\ \mathrm{fm}$ to $2\ \mathrm{fm}$. Utilizing these models, NM properties can be calculated using the relativistic mean field (RMF) approach Walecka (1974); Wiringa et al. (1988); Gao et al. (2022); Gao and Harada (2024), which is the most practical and economical framework to introduce density effects.
@@
It is recognized that Walecka-type models lack the consideration of QCD symmetry patterns, the valid region of effective operators, and theoretical errors. Chiral effective field theory ($\rm{\chi EFT}$), thanks to the pioneering works by Weinberg Weinberg (1979, 1991), offers a powerful framework for studying nuclear forces at long ranges, anchored on QCD symmetry. To develop a realistic model of nuclear forces, it is accepted that vector mesons—$\rho$ and $\omega$— and isoscalar-scalar meson $\sigma$ are indispensable. The former can be regarded as gauge fields of hidden local symmetry (HLS) in $\rm{\chi EFT}$  Bando et al. (1985, 1988); Harada and Yamawaki (2003). However, the inclusion of the $\sigma$ meson as an independent degree of freedom in $\rm{\chi EFT}$ is not straightforward, since the fourth scalar component of the chiral four-vector is integrated out when transitioning from the linear sigma model to the nonlinear sigma model—the leading term of $\rm{\chi EFT}$. Considering that the lowest-lying scalar meson $\sigma$ has a mass roughly equivalent to that of the kaon, which is a Nambu-Goldstone (NG) boson of chiral symmetry breaking in three-flavor $\rm{\chi EFT}$ ($\chi\rm EFT_{3}$), by supposing that QCD has a nonperturbative infrared fixed point, Crewther and Tunstall Crewther and Tunstall (2015); Catà et al. (2019) proposed that the lowest-lying scalar meson can be considered as the NG boson (dilaton) of the spontaneous breaking of scale symmetry. Consequently, $\rm{\chi EFT}$ is extended to include the scalar meson, resulting in the scale-chiral effective theory ($\chi\rm EFT_{\sigma}$).
@@
Based on the terminologies of $\chi\rm EFT_{\sigma}$ and HLS, the $\chi\rm EFT$ with baryons is constructed in Refs. Lee et al. (2015); Li et al. (2017) at the leading chiral order to discuss nucleon interactions. It is denoted as bsHLS, with ‘b’ for baryon and ‘s’ for scale. In bsHLS, the potential from meson exchanges in OBE models can be reproduced by expanding bsHLS to the first order of linear field couplings but with additional symmetry considerations. By using the low-momentum potential $V_{\rm lowk}$ renormalization group approach Bogner et al. (2003); Holt et al. (2007), taking the “leading order scale symmetry (LOSS)” approximation where the trace anomaly effect enters only through the dilaton potential, which breaks scale symmetry explicitly and spontaneously, we found that the chiral-scale EFT with few parameters can successfully describe not only NM at the saturation density but also the compact-star matter at $n=(5-7)n_{0}$ Ma and Rho (2020a). A novel phenomena, which was not realized before, is that in nuclear matter at densities $n=(2-4)n_{0}$ the sound velocity of NM saturates the conformal limit $v_{s}^{2}=1/3$, but the trace of the energy-momentum tensor does not vanish, that is, the NM exhibits a pseudo-conformal structure Ma and Rho (2019a, b); Ma et al. (2019). And the corrections to LOSS in the baryonic part have been found crucial for understanding the quenched $g_{A}$ value in the super-allowed Gamow-Teller transitions of heavy nuclei Lubos et al. (2019); Ma and Rho (2020b).
@@
In this work, we study NM properties using bsHLS with the RMF method. After identifying the NM properties around $n_{0}$, it was found that bsHLS using the RMF method can yield reasonable results. The data can be reproduced with the choice of $\beta^{\prime}=\frac{\partial\beta}{\partial\alpha}\sim 1$, consistent with the results of Refs. Ma and Rho (2018); Shao and Ma (2022), where a dilaton limit fixed point is assumed in the medium. The combination of $f_{\chi}m_{\sigma}$ is constrained to be $\simeq 2.3\times 10^{5}$ MeV2 which is consistent with what estimated before in the skyrmion crystal approach such that physically interesting results can be obtained Lee et al. (2003); Park et al. (2008). By comparing the symmetry energy $E_{\rm sym}$ and incompressibility coefficient $K(n)$ obtained from our bsHLS and Walecka-type models, we find that the bsHLS can make the symmetry energy stiff at subsaturation density but soft at intermediate densities to meet the constraints of GW170817 Abbott et al. (2018) and the neutron skin thickness of ${}^{208}\rm Pb$ Reed et al. (2021) simultaneously, without introducing additional freedoms such as the $\delta$ meson in Refs. Li et al. (2022); Ma and Ma (2024). In addition, we find that the incompressibility coefficient of bsHLS surges at intermediate densities, while the Walecka-type models exhibit gentler behaviors, resulting in a better description of NS structures using bsHLS. The maximum mass of NS can nearly reach $3M_{\odot}$ with a pure hadron phase, meeting the constraints of GW170817 Abbott et al. (2018, 2019) and PSR J0740+6620 Miller et al. (2021); Riley et al. (2021), whereas Walecka-type models can only reach $2M_{\odot}$ within these constraints, as analyzed in Ref. Guo et al. (2024). We find that this is due to the kink behavior of $\sigma$ expectations at intermediate densities, induced by the nonlinear realization of scale symmetry. Moreover, the value of $\beta^{\prime}$, behavior of $\langle\chi\rangle^{*}$ and Brown-Rho scaling (B-R scaling) at different densities can significantly affect NS structures, indicating a relationship between QCD symmetry patterns and macroscopic phenomena.
@@
In this section, we establish the theoretical framework of bsHLS and derive the equation of state (EoS) for NM within the relativistic mean field (RMF) approach. Focusing on NM only composed of nucleons, we restrict the bsHLS Lagrangian to the two-flavor case, involving only $u$ and $d$ quarks. The leading order bsHLS Lagrangian $\mathcal{L}$ can be decomposed into the baryonic part $\mathcal{L}_{B}$ and the mesonic part $\mathcal{L}_{M}$:
@@
$g_{\omega NN}=\frac{1}{2}\left(g_{V_{p}}C_{V_{p}}+g_{V_{0}}C_{V_{0}}-1\right)g_{\omega}\ ,$
$g_{\rho NN}=\frac{1}{2}\left(g_{V_{p}}C_{V_{p}}-1\right)g_{\rho}\ ,$
@@ III Phenomenological analysis
In the phenomenological analysis, we choose vacuum values $f_{\pi}=92.4~{}\rm MeV$, $m_{N}=939~{}\rm MeV$, $m_{\pi}=140~{}\rm MeV$, $m_{\omega}=783~{}\rm MeV$ and $m_{\rho}=765~{}\rm MeV$ Workman et al. (2022). The free parameters are $M_{\sigma}=m_{\sigma}f_{\chi}$, $\beta^{\prime}$, $r$, $g_{\omega NN}$, $g_{\rho NN}$, $g^{SSB}_{\omega NN}$ and $g^{SSB}_{\rho NN}$, which can be estimated by the properties of NM around saturation density $n_{0}$. The calculated results of NM properties are listed in Table 1, and the corresponding low energy constants (LECs) are given in Table 2. It can be seen that NM properties obtained from both bsHLS-L and bsHLS-H are consistent with empirical values.
@@ III.1 Nuclear matter properties
The properties of NM is analyzed at first. From Table 1, one can see that, with appropriate choices of parameters in Table 2, our bsHLS can yield NM properties around (sub)saturation density, satisfying the constraints from empirical data.
@@ III.2 Neutron star structures
Next, the NS structures are studied using the EOSs discussed above for pure neutron matter (PNM). The results of NS structures are shown in Fig. 2 and Table 4.
Figure 2: The M-R relations from bsHLS and Walecka-type models. The constraints are estimated from Refs. Abbott et al. (2018, 2019); Riley et al. (2019, 2021). The M-R relation is calculated by solving the Tolman-Oppenheimer-Volkoff (TOV) equation Oppenheimer and Volkoff (1939); Tolman (1939). All EoSs for TOV equation calculation are interpolated to a BPS EoS Baym et al. (1971) below $0.01\ \rm fm^{-3}$ Arnett and Bowers (1977).
Table 4: Tidal deformations from bsHLS and Walecka-type models defined in Eq. (14) and Table 3 for NS with mass of $1.4\ M_{\odot}$. They are calculated from the EOSs used for the M-R relation in Fig. 2, using the formalisms in Ref. Postnikov et al. (2010).
@@
From Fig. 2, one can see that the mass-radius (M-R) relation obtained from bsHLS-H falls within the constraints of astro-observations. Although the result from FSU-$\delta 6.7$ satisfies these constraints, it requires introducing a new degree of freedom $\delta$. The Walecka-type models without $\delta$ considered in this work cannot provide results within the constraints of NSs, though they reproduce the NM properties around $n_{0}$ Horowitz and Serot (1981); Reinhard et al. (1986); Sugahara and Toki (1994). This highlights the advantage of bsHLS in interpreting the structures of nuclei and NSs within a unified framework. It should be emphasized that the present analysis indicates that, in order to have realistic parameter space of nuclear model, including Walecka-type models, the data of both NM properties and NS structures should be considered in the statistical analysis simultaneously, as discussed in Ref. Guo et al. (2024). When meeting the NS constraints, the $M_{\rm max}$ of NSs predicted by these Walecka-type models and chiral nuclear force models is around or slightly above $2M_{\odot}$, as discussed in Refs. Özel and Freire (2016); Guo et al. (2024); Fan et al. (2024). This makes bsHLS more appealing, as the predicted $M_{\rm max}$ of bsHLS-H is $\sim 2.8M_{\odot}$. This predicted $M_{\rm max}$ may have profound implications for addressing the gap problem in the continuous mass distribution of supernova remnants Bailyn et al. (1998); Farr et al. (2011) and for understanding the gravitational wave event GW190814 Abbott et al. (2020).
@@
In summary, the above discussion indicates that bsHLS-H is a reasonable model for NM and NS. It implies that the scaling parameter $r\approx 0.19$, $\beta^{\prime}\approx 1.15$, $M_{\sigma}=m_{\sigma}f_{\chi}\approx 2.3\times 10^{5}~{}\rm MeV^{2}$, and the couplings between vector mesons and nucleons $g_{\rho NN}\approx 4.17$, $g_{\omega NN}\approx 11.0$. The scaling parameter $r\approx 0.19$ is consistent with pion-nuclei bound state data Kienle and Yamazaki (2004). And the $\beta^{\prime}\approx 1.15$ agrees with estimates from skyrmion crystal approach Ma and Rho (2018); Shao and Ma (2022). If $f_{\chi}$ is taken to be $3f_{\pi}\approx 270~{}\rm MeV$, $m_{\sigma}\approx 850~{}\rm MeV$, while the power-counting mechanism of Crewther and Tuntall remains valid Crewther and Tunstall (2015). The coupling constant $g_{\rho NN}\approx 4.17$ aligns with the results from the OBE potential analysis of nucleon interactions Erkelenz (1974), and $g_{\omega NN}\approx 11.0$ is in agreement with the analyses of nucleon-nucleon scatterings Backman et al. (1985).
@@ III.3 The patterns of scale symmetry and phenomenologies
The difference between bsHLS-L and bsHLS-H becomes significant in the M-R relation, as shown in Fig. 2, though it is not distinguishable from the NM properties around saturation density. The key difference lies in $\beta^{\prime}$, which is connected to the existence of the pseudo-conformal structure at high densities Shao and Ma (2022). The $\beta^{\prime}$ of bsHLS-H aligns with the constraints of the pseudo-conformal limit, whereas the $\beta^{\prime}$ of bsHLS-L does not. This highlights the impact of scale symmetry patterns on NS structures.
@@
And, as shown in Fig. 3, the M-R relations for PNM are roughly consistent with observational constraints, but there is a decrease of $M_{\rm max}$ compared to bsHLS-H.
@@
Additionally, the NM properties are investigated with B-R scaling turned off. We find that with the parameter set bsHLS-N: $M_{\sigma}=1.11\times 10^{5}\ \rm MeV^{2}$, $\beta^{\prime}=1.10$, $g_{\omega NN}=-11.0$, $g_{\rho NN}=4.17$, $g^{SSB}_{\omega NN}=-192$ and $g^{SSB}_{\rho NN}=4.85$, the NM properties can still be reproduced (e.g., $n_{0}=0.16\ \rm{fm}^{-3}$, $e_{0}=-16.0\ \rm{MeV}$, $K_{0}=241\ \rm{MeV}$, $E_{{\rm sym}}(n_{0})=29.3\ \rm{MeV}$). However, this approach results in a much softer scale symmetry parameter $\langle\chi\rangle^{*}$ flow compared to the cases with B-R scaling. As a result, the M-R relation becomes unnatural and deviates from the constraints, as shown in Fig. 5.